a ball is dropped from a height h from rest.the ball travels (h/2) in last 1.0s the total time taken to the ball hit the ground is...
Ans 3.41s
1
Expert's answer
2020-09-22T15:36:32-0400

"t-t_0=1\to \sqrt{\frac{2h}{g}}-\sqrt{\frac{2h/2}{g}}=1\to"


"\sqrt h(\sqrt{\frac{2}{g}}-\sqrt{\frac{1}{g}})=1\to"


"h=\frac{1}{(\sqrt{\frac{2}{g}}-\sqrt{\frac{1}{g}})^2}=57.197(m)"


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\cdot 57.197}{9.81}}=3.41(s)"